The fifth number plus the third number equals fourteen.
The fourth number is one more than the second number.
The first number is one less than twice the second number.
The second number plus the third number equals ten.
The sum of all five numbers is 30.


Answer by stanbon(75887)

The fifth number plus the third number equals fourteen.
The fourth number is one more than the second number.
The first number is one less than twice the second number.
The second number plus the third number equals ten.
The sum of all five numbers is 30.
--------------------------------------------
Let the numbers, in order, be a,b,c,d,e,f
EQUATIONS:
a +b + c +d +e +f = 30
0 +b + c = 10
a +2b = -1
-b +d = 1
c +e = 14
--------
Comment: You need one more equation statement.
======================
Cheers,
Stan H.